# Polyrating: A Cost-Effective and Bias-Aware Rating System for LLM Evaluation

### v2

@@ 3.2 Polyrating
However, while traditional benchmarks often predict a ranking similar to human preference data, their absolute performance measurements do not always align well with human ratings. For example, an easy benchmark will cluster all models close together because weaker models can answer most questions correctly, making it impossible for stronger models to achieve very high win rates against them. This makes Polyrating less effective, since the ratings from human preference data are spread further apart than those from these easy benchmarks. To address this issue, we introduce a hyperparameter that rescales the observed win rates of traditional benchmarks. This hyperparameter is optimized on the training data to realign Polyrating with the human preference data, allowing for a more effective combination of the two. Full details on this parameter are in App.˜D.
@@
In App.˜C, we show that the optimization objective from Eq.˜3 is convex and twice differentiable with respect to $\mathbf{R}_{\text{base}},\bm{\alpha}$ and $\bm{\beta}$, allowing us to use standard optimization techniques to optimize the loss with respect to these parameters. Specifically, we use Newton’s method for the model-specific parameters and L-BFGS for the shared parameters. Furthermore, we prove in App.˜C that, under weak assumptions, the obtained ratings converge to the ratings that maximize predictive capabilities and thus fulfill the final design goal.
@@
We compute pivotal intervals using bootstrapping to obtain rating uncertainties (Tibshirani, 1984). Specifically, given the actual estimate of the ratings $\hat{R}$ and $n$ bootstrap estimates $\hat{R}_{1},\dots,\hat{R}_{n}$, the pivotal interval for a confidence $\alpha$ is $[\hat{R}-\hat{R}_{(1-\alpha/2)},\hat{R}-\hat{R}_{(\alpha/2)}]$, where $\hat{R}_{(1-\alpha/2)}$ and $\hat{R}_{(\alpha/2)}$ are respectively the $1-\alpha/2$ and $\alpha/2$ quantiles of the bootstrap estimates. For brevity, we report the $2\sigma$ confidence intervals obtained using bootstrapping in Section˜4, with detailed pivotal intervals reported in App.˜E.
@@ 4.1 Bias Detection
We briefly explain the measured biases and refer to App.˜D for a full overview of all biases along with their functional form. We include the well-known length bias (Dubois et al., 2024; Singhal et al., 2023), which measures bias with respect to the length of the completion, and position bias (Shi et al., 2024; Wang et al., 2023), which measures bias with respect to the order of the models in the game. We further use classifiers (Babakov et al., 2023; Camacho-collados et al., 2022) to evaluate the influence of formality and sentiment of the model’s output. Finally, we compute the repetitiveness of the answer by measuring the number of unique tokens in the completion and check the influence of the readability of an answer using the Flesch Reading Ease score (Kincaid et al., 1975). We model all of these biases as shared features when fitting Polyrating, allowing us to accurately estimate their influence on the resulting ratings.
Table 1: Fitted coefficients for the biases and their average influence on model ratings for both human and LLM-based evaluation. The functional form of $f_{\text{bias}}$ used for each bias can be found in App.˜D. The influence is computed as $\mathbb{E}_{g}(\alpha_{\text{bias}}\cdot|f_{\text{bias}}(g,0)-f_{\text{bias}}(g,1)|)$.
@@ 4.2 Improved Sample Efficiency
We demonstrate that Polyrating is more sample-efficient compared to traditional univariate approaches and therefore reduces the cost of human evaluation.
@@
We use the full Chatbot Arena dataset (Chiang et al., 2024b), which contains over one million questions across various tasks. Each question is answered by two models and judged by the human that posed the question. We use the tasks that contain Chinese, code-based, and hard questions for this experiment and refer for a full description of these tasks to App.˜D.
@@
Figure 3: Comparison between Polyrating and the univariate baseline when leveraging information from existing benchmarks. For the left and middle plot, the $x$-axis shows the number of human annotations used. For the right plot, the $x$-axis shows the amount of samples from the Chinese code task. The logistic loss is normalized by subtracting the loss of the best possible rating. The grey horizontal line indicates the loss of a rating system that assigns the same rating to all models.
@@ Results
Table˜2 and Table˜3 show the ratings of several models in the leaderboard fitted using Polyrating and a univariate approach respectively. For a full overview of all models, we refer to App.˜E. By examining the modifiers computed by Polyrating, we immediately see the downside of the univariate approach. In both the English and Chinese tasks, the Mixtral model performs very differently compared to the other tasks. Fixing the shifting constant using this model results in significant ratings shifts for these two tasks compared to the base rating.
This effect is most apparent in the Chinese task. Almost all models gain substantial ratings in the univariate approach, even though yi-large-preview is the only model built bilingually from the ground up. In contrast, Polyrating shows that yi-large-preview is the only model that gains rating in this task. Similarly, llama-3-70b-instruct, a model specifically trained for English, gains only 22 rating points in the English task using the univariate approach, whereas Polyrating shows it gains 67 rating points, aligning more with expectations. The last two tasks show more similarities between the two approaches, but we note that the arbitrary shifting constant for each task makes it very hard to actually trust these modifiers in the univariate approach.
Table 2: Several models in a multidimensional leaderboard fitted using Polyrating. The given rank of the models indicates its rank in the complete leaderboard shown in LABEL:tab:leaderboard in App.˜E.
@@
Table 3: Several models in the leaderboard fitted using a unidimensional approach where each task is fitted separately. Modifiers for the tasks are shown instead of the fitted rating to make comparison with Polyrating easier. The complete leaderboard is shown in LABEL:tab:leaderboard-single in App.˜E.
@@ 5 Related Work
Preference datasets for LLMs have become common to evaluate model capabilities in areas lacking ground-truth benchmarks. The most popular one is the Chatbot Arena (Chiang et al., 2024b), which contains over one million user queries and evaluates models in various tasks such as code, math, and multilingual understanding. Wildbench (Lin et al., 2024), MT-Bench (Zheng et al., 2023), and AlpacaEval (Dubois et al., 2024) are LLM-based evaluation frameworks that have gained attention. Among these, AlpacaEval is the only one that applies a length-control bias similar to Polyrating to obtain a higher correlation with human judges. However, this fitted bias is not directly interpretable and is less generic than the approach used in Polyrating. Additionally, Polyrating employs priors on various terms to improve sample efficiency and eliminate shift-invariance, which AlpacaEval lacks.
@@
Multivariate rating systems have been used before in recommendation systems (Chen et al., 2018; Adomavicius et al., 2005; Kong et al., 2019). These developed systems are extensions to the more classical Elo (Elo, 2008) and Glicko (Glickman, 2002) rating systems. However, they are not directly applicable to LLM evaluation, as they do not take into account the specific biases and dependencies that are present in LLM evaluation. Furthermore, the limited numbers of models allow us to build an exact optimization algorithm, unlike in recommendation systems where approximate algorithms are necessary due to the high number of products.
@@ 7 Conclusion
We have included our code in the supplementary material with instructions how to run and reproduce all the results presented in the paper. Furthermore, App.˜C contains detailed proofs of all theoretical statements made in the paper, particularly with respect to the optimality of Polyrating.
Acknowledgements
This work was funded in part by the Swiss National Science Foundation (SNSF) [200021_207967].
This work has been done as part of the EU grant ELSA (European Lighthouse on Secure and Safe AI, grant agreement no. 101070617). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or European Commission. Neither the European Union nor the European Commission can be held responsible for them.
The work has received funding from the Swiss State Secretariat for Education, Research and Innovation (SERI).
@@ Appendix B Alternative Rating Systems
Appendix C Proofs
@@
where the first inequality follows from Eq.˜9 and the last from Eq.˜7 and Eq.˜8. Since Eq.˜10 ensures that the difference in the bias term can at most differ by $\epsilon/4$, we find $\mathcal{L}_{\text{full}}^{(n)}(D^{(n)},\mathbf{R^{\prime\prime}}^{(n)})<\mathcal{L}_{\text{full}}^{(n)}(D^{(n)},\mathbf{R^{\prime}}^{(n)})$. Therefore, $\mathbf{R^{\prime}}^{(n)}_{\text{task}}$ cannot be the optimal solution, which is a contradiction to the optimality of $\mathbf{R^{\prime}}^{(n)}$. Therefore, $\mathbf{R^{\prime}}_{\text{task}}$ must converge to $\mathbf{R}_{\text{task}}^{*}$.
Appendix D Experimental Details
In this section, we provide detailed descriptions of the biases and tasks used in our experiments. In Table˜4 we describe the biases used in Section˜4.1 in more detail. In Table˜5 we describe the tasks used in Section˜4 in more detail.
@@ Appendix D Experimental Details
Appendix E Detailed Results

### v3

@@ 3.2 Polyrating
However, while traditional benchmarks often predict a ranking similar to human preference data, their absolute performance measurements do not always align well with human ratings. For example, an easy benchmark will cluster all models close together because weaker models can answer most questions correctly, making it impossible for stronger models to achieve very high win rates against them. This makes Polyrating less effective, since the ratings from human preference data are spread further apart than those from these easy benchmarks. To address this issue, we introduce a hyperparameter that rescales the observed win rates of traditional benchmarks. This hyperparameter is optimized on the training data to realign Polyrating with the human preference data, allowing for a more effective combination of the two. Full details on this parameter are in App.˜E.
@@
In App.˜D, we show that the optimization objective from Eq.˜3 is convex and twice differentiable with respect to $\mathbf{R}_{\text{base}},\bm{\alpha}$ and $\bm{\beta}$, allowing us to use standard optimization techniques to optimize the loss with respect to these parameters. Specifically, we use Newton’s method for the model-specific parameters and L-BFGS for the shared parameters. Furthermore, we prove in App.˜D that, under weak assumptions, the obtained ratings converge to the ratings that maximize predictive capabilities and thus fulfill the final design goal.
@@
We compute pivotal intervals using bootstrapping to obtain rating uncertainties (Tibshirani, 1984). Specifically, given the actual estimate of the ratings $\hat{R}$ and $n$ bootstrap estimates $\hat{R}_{1},\dots,\hat{R}_{n}$, the pivotal interval for a confidence $\alpha$ is $[\hat{R}-\hat{R}_{(1-\alpha/2)},\hat{R}-\hat{R}_{(\alpha/2)}]$, where $\hat{R}_{(1-\alpha/2)}$ and $\hat{R}_{(\alpha/2)}$ are respectively the $1-\alpha/2$ and $\alpha/2$ quantiles of the bootstrap estimates. For brevity, we report the $2\sigma$ confidence intervals obtained using bootstrapping in Section˜4, with detailed pivotal intervals reported in App.˜F.
@@ 4.1 Bias Detection
We briefly explain the measured biases and refer to App.˜E for a full overview of all biases along with their functional form. We include the well-known length bias (Dubois et al., 2024; Singhal et al., 2023), which measures bias with respect to the length of the completion, and position bias (Shi et al., 2024; Wang et al., 2023), which measures bias with respect to the order of the models in the game. We further use classifiers (Babakov et al., 2023; Camacho-collados et al., 2022) to evaluate the influence of formality and sentiment of the model’s output. Finally, we compute the repetitiveness of the answer by measuring the number of unique tokens in the completion and check the influence of the readability of an answer using the Flesch Reading Ease score (Kincaid et al., 1975). We model all of these biases as shared features when fitting Polyrating, allowing us to accurately estimate their influence on the resulting ratings.
Table 1: Fitted coefficients for the biases and their average influence on model ratings for both human and LLM-based evaluation. The functional form of $f_{\text{bias}}$ used for each bias can be found in App.˜E. The influence is computed as $\mathbb{E}_{g}(\alpha_{\text{bias}}\cdot|f_{\text{bias}}(g,0)-f_{\text{bias}}(g,1)|)$.
@@ 4.2 Improved Sample Efficiency
We demonstrate that Polyrating is more sample-efficient compared to traditional univariate approaches and therefore reduces the cost of human evaluation. In App.˜C, we show that a slight adjustment of Polyrating additionally enables an improved sample efficiency when evaluating new model versions.
@@
We use the full Chatbot Arena dataset (Chiang et al., 2024b), which contains over one million questions across various tasks. Each question is answered by two models and judged by the human that posed the question. We use the tasks that contain Chinese, code-based, and hard questions for this experiment and refer for a full description of these tasks to App.˜E.
@@
Figure 3: Comparison between Polyrating and the univariate baseline when leveraging information from existing benchmarks. For the left and middle plot, the $x$-axis shows the number of human annotations used. For the right plot, the $x$-axis shows the amount of samples from the Chinese code task. The logistic loss is normalized by subtracting the loss of the best possible rating.
@@ Results
Table˜2 and Table˜3 show the ratings of several models in the leaderboard fitted using Polyrating and a univariate approach respectively. For a full overview of all models, we refer to App.˜F. By examining the modifiers computed by Polyrating, we immediately see the downside of the univariate approach. The Mixtral model performs significantly worse, resp. better, on the Chinese, resp. English, task compared to its base rating. Therefore, fixing the shifting constant using Mixtral results in significant ratings shifts for these tasks making cross-task comparisons impossible.
This effect is most apparent in the Chinese task. 97 of the 114 models included in the dataset gain rating for this task in the univariate approach, even though most models were not specifically trained for the Chinese language. In contrast, Polyrating shows that only 67 models gain rating, and that the only ones to do so significantly are models trained by Chinese model providers, such as Yi, Qwen and GLM.
Inspecting the English task, we find that Polyrating indicates that top models tend to lose rating for this task, while bad models tend to gain rating. This is in line with expectations as older and worse models were predominantly trained on English data, making them more suitable for this task and thus increase their rating. However, the univariate approach simply shows that more than 100 models lose rating for this task, with no discernable pattern.
Table 2: Several models in a multidimensional leaderboard fitted using Polyrating. The given rank of the models indicates its rank in the complete leaderboard shown in LABEL:tab:leaderboard in App.˜F.
@@
Table 3: Several models in the leaderboard fitted using a unidimensional approach where each task is fitted separately. Modifiers for the tasks are shown instead of the fitted rating to make comparison with Polyrating easier. The complete leaderboard is shown in LABEL:tab:leaderboard-single in App.˜F.
@@ 5 Related Work
Preference datasets for LLMs have become common to evaluate model capabilities in areas lacking ground-truth benchmarks. The most popular one is the Chatbot Arena (Chiang et al., 2024b), which contains over one million user queries and evaluates models in various tasks such as code, math, and multilingual understanding. Wildbench (Lin et al., 2024), MT-Bench (Zheng et al., 2023), and AlpacaEval (Dubois et al., 2024) are LLM-based evaluation frameworks that have gained attention. Among these, AlpacaEval is the only one that applies a length-control bias similar to Polyrating to obtain a higher correlation with human judges. However, this fitted bias is not directly interpretable and is less generic than the approach used in Polyrating. Additionally, Polyrating employs priors on various terms to improve sample efficiency and eliminate shift-invariance, which AlpacaEval lacks. Therefore, AlpacaEval cannot be used to obtain any of the benefits of Polyrating that were presented in Section˜4.
@@
Multivariate rating systems have been used before in recommendation systems (Chen et al., 2018; Adomavicius et al., 2005; Kong et al., 2019; Abdi et al., 2021). These developed systems are extensions to the more classical Elo (Elo, 2008) and Glicko (Glickman, 2002) rating systems. However, they are not directly applicable to LLM evaluation, as they do not take into account the specific biases and dependencies that are present in LLM evaluation. Furthermore, the limited numbers of models allow us to build an exact optimization algorithm, unlike in recommendation systems where approximate algorithms are necessary due to the high number of rated players (or products). These approximate algorithms are not suitable for LLM evaluation, as shown in App.˜B. Furthermore, these systems do not include priors on the ratings, which are crucial for the sample efficiency of Polyrating.
@@ 7 Conclusion
We have included our code in the supplementary material with instructions how to run and reproduce all the results presented in the paper. Furthermore, App.˜D contains detailed proofs of all theoretical statements made in the paper, particularly with respect to the optimality of Polyrating.
@@ Appendix B Alternative Rating Systems
First of all, we see that the approximate Glicko system [Glickman, 2002] performs by far the worst, as expected. Using approximate systems for LLM evaluation is not recommended, as these systems were designed for time-varying, large-scale rating systems for multiple million players.
@@
Appendix C Model Versions
Models are iteratively improved through time. Evaluating these iterations is essential for tracking performance changes and understanding the impact of updates. For developers who create multiple versions of a model simultaneously or in quick succession, it is essential to evaluate the relative performance of these versions as cost-effectively as possible. However, this process can be computationally expensive, especially when human evaluations are used. This section describes how Polyrating enables more efficient evaluation of model versions.
C.1 Incorporating Model Versions
Model versions evolve over time, much like human capabilities in competitive games. Rating systems often reflect this evolution by introducing time-dependent factors, such as the method proposed by Coulom [2008], which uses a time-dependent Bradley-Terry (BT) model. This approach incorporates a Gaussian prior on consecutive ratings, ensuring they do not shift arbitrarily over time:
$R_{t+1}-R_{t}\sim\mathcal{N}(0,\sigma^{2})$
Unlike human capabilities, model updates are irregular. Some models may never be updated, while others evolve rapidly. However, Coulom [2008] assumes that the ratings are updated at regular intervals. To remove this assumption, we modify the prior by only updating ratings upon a new release. For two versions $v_{1}$ and $v_{2}$, we impose a regularizing prior on the base rating:
$R^{v_{2}}_{\text{base}}-R^{v_{1}}_{\text{base}}\sim\mathcal{N}(0,\sigma^{2})$
This adjustment enables faster convergence by constraining rating shifts between versions without requiring regularity. Using this prior, the loss function for evaluating versions becomes:
$\mathcal{L}_{\text{version}}(D,\mathbf{R}_{\text{base}},\bm{\alpha},\bm{\beta})=\mathcal{L}_{\text{full}}(D,\mathbf{R}_{\text{base}},\bm{\alpha},\bm{\beta})+\sum_{(v_{1},v_{2})\in V_{\mathcal{M}}}\frac{(R^{v_{1}}_{\text{base}}-R^{v_{2}}_{\text{base}})^{2}}{2\sigma^{2}}$ (4)
Here, $V_{\mathcal{M}}$ represents all pairs of consecutive model versions. Thus, if a model has versions $v_{1},\dots,v_{n}$, then $\{(v_{1},v_{2}),\dots,(v_{n-1},v_{n})\}\subset V_{\mathcal{M}}$. Note that we only consider versions of minor updates. For major updates (e.g., GPT-3.5 to GPT-4), the update is not included in $V_{\mathcal{M}}$ as the rating of the new model will likely be substantially higher and is not related to the previous model anymore. This loss function can be optimized using the same techniques as described in Section˜3. Therefore, no additional adjustments are needed to evaluate model versions using Polyrating.
Figure 5: Convergence rate of the univariate method and Polyrating when evaluating model versions. The x-axis represents the number of games available in the training set associated with the subsequent versions, while the y-axis represents the loss of the univariate method and Polyrating.
C.2 Evaluating Model Versions
We evaluate whether Polyrating can effectively reduce the evaluation costs associated with assessing new model versions. To do this, we adopt an experimental setup similar to the one described in Section˜4. Using the Chatbot Arena dataset, we first identify models with multiple versions to construct the set $V_{\mathcal{M}}$. Out of the 129 models in the dataset, 43 are subsequent versions of earlier models. We then split the dataset into two parts: one containing games involving these subsequent versions, and the other including all remaining games. As before, the latter set is used for both the univariate method and Polyrating, while the former set is further divided into training and test splits.
[…]
Appendix D Proofs
@@
where the first inequality follows from Eq.˜10 and the last from Eq.˜8 and Eq.˜9. Since Eq.˜11 ensures that the difference in the bias term can at most differ by $\epsilon/4$, we find $\mathcal{L}_{\text{full}}^{(n)}(D^{(n)},\mathbf{R^{\prime\prime}}^{(n)})<\mathcal{L}_{\text{full}}^{(n)}(D^{(n)},\mathbf{R^{\prime}}^{(n)})$. Therefore, $\mathbf{R^{\prime}}^{(n)}_{\text{task}}$ cannot be the optimal solution, which is a contradiction to the optimality of $\mathbf{R^{\prime}}^{(n)}$. Therefore, $\mathbf{R^{\prime}}_{\text{task}}$ must converge to $\mathbf{R}_{\text{task}}^{*}$.
Appendix E Experimental Details
In this section, we provide detailed descriptions of the biases and tasks used in our experiments. In Table˜4 we describe the biases used in Section˜4.1 in more detail. In Table˜5 we describe the tasks used in Section˜4 in more detail. Finally, we note that any run using Polyrating took at most 6 hours on a single CPU, even for huge datasets with a million samples, 100 models and 10 tasks.
@@ Appendix E Experimental Details
Appendix F Detailed Results